# Electronic structure and conductivity in functionalized multi-layer Black Phosphorene

### v1

@@ III Numerical method for transport
where $S$ is the surface and $E_{F}^{0}$ is the Fermi energy of the system without static defect, thus $E_{F}^{0}$ is at the gap middle of the gap. Note that in the presence of defect band states in the gap that has a non-zero $\sigma_{M}$, the DOS at $E_{F}$ can be non-zero, and therefore the mobility is not well defined because the expression (8) diverges ($N_{e}$ goes to zero for $\mu_{c}=E_{F}$). This is because the mobility is defined for a semiconductor but not for systems with states at $E_{F}$.
IV Results and discussion
In this section, we present the results on the effect of vacancies (resonant scatterers) and Anderson disorder (non-resonant scatterers) on the electronic properties of monolayer BP, few-layer BP, and bulk BP.
@@ IV.1 Resonant scatterers
FIG. 3a illustrates the density of states (DOS) for a vacancy concentration of $c=1\%$ (resonant scatterers) relative to the total number of atoms in monolayer, few-layer BP, and bulk BP. Notably, there is a decrease in the gap as the number of layers increases. The FIG. 3b shows that overall, for few-layer BP (FIG. 3d for 3 layers) and bulk configurations, the gap increases as the vacancy concentration rises, except for the monolayer case where the gap remains nearly constant (FIG. 3c).
@@
We can clearly distinguish the midgap state peak due to vacancies in the gap (FIG. 3a). It is not centered at zero energy because the coupling terms beyond the first neighbors break the electron-hole symmetry of the bipartite Hamiltonian. Its width remains almost constant as the number of layers increases showing that its energy dispersion is mainly due to intralayer couplings between midgap states. As the gap decreases with an increasing number of layers, the specific value of $c$ at which the midgap peaks completely fill the gap depends on the number of layers. For monolayer, at studied $c$ concentration, the gap is not filled and we can distinguish that secondary peaks appear (see FIGs. S1 in SM [32]). On the other hand, the gap is completely filled for $c>1-2\%$ in the case of 5 layers and a bulk, respectively (see FIGs. S2 and S3 in SM [32]). In FIG. 3b, the points corresponding to these particular cases are not represented.
@@
FIGs. 3e,f show the microscopic conductivity multiplied by concentration versus energy for monolayer and trilayer according to the $x$-direction $\sigma_{xx}$ (bold line) –armchair direction– and according to the $y$-direction $\sigma_{yy}$ (dashed line) –zigzag direction–. See also FIG. S4 in SM [32] for bulk BP. The armchair conductivity $\sigma_{xx}$ is greater than the zigzag conductivity $\sigma_{yy}$ for all cases (monolayer, few-layer and bulk). The proportionality between $\sigma_{xx}$ and $\sigma_{yy}$ remains the same for all systems. This ratio is a signature of the structural anisotropy of these materials, which results in an anisotropy of the effective masses of electrons and holes. [34, 69].
[…]
@@
We now study the effect of vacancies on the conductivity and mobility at room temperature ($T=300$ K) versus the charge carriers $N_{e}$ (Eqs. (6), (8) and (9)). The prediction of other works are not similar and depends roughly with the numerical study [31]. FIG. 4a shows the armchair conductivity per layer versus $N_{e}$ for monolayer, few-layer BP, and bulk with a concentration of vacancies of $c=1\%$. The conductivity of holes is larger than that of electrons for high values of charge carriers. The zigzag conductivity presents the same behavior (see FIG. S5 in SM Ref. [32]). These results are in agreement with the experimental results found by Xia et al. [7] for a BP thin film.
@@
To compare the effect of the concentration of vacancies, we present the armchair conductivity versus $N_{e}$ for monolayer (bold line) and trilayer (dashed line) with different concentrations that are shown in FIG. 4b. As expected, the conductivity decreases when $c$ increases, moreover we can remark a relative asymmetry concerning the nature of charge carriers (electrons or holes). In FIG. 4c, the ratio $\sigma_{xx}/\sigma_{yy}$ is shown for monolayer, few-layer BP and a bulk versus $N_{e}$. $\sigma_{xx}$ is more larger than $\sigma_{yy}$ regardless of charge carrier nature. This result is the same than the theoretic predictions [10, 37] and with experimental results [38]. It is easy to see that the anisotropy in the behavior of conductivity is more significant for electrons than for holes for all systems. However, the ratio $\sigma_{xx}/\sigma_{yy}$ is larger for monolayer than for the other systems for holes whereas monolayer ratio is smaller for small electron density. This ratio is also independent of a number of vacancies as found previously (Boltzmann transport). This is shown in FIG. 4d for monolayer and trilayer, and this is valid for all systems. Note that the ratio $\sigma_{xx}/\sigma_{yy}$ is equal to the mobility ratio $\mu_{xx}/\mu_{yy}$.
@@
To enable simpler comparisons with experimental results, we plot the mobility at room temperature versus $N_{e}$ (FIG. 4e), according to formulas (8) and (9), for monolayer, few layers and the bulk for $c=1\%$. In the case of the monolayer, the mobility is of the same order of magnitude regardless of the charge carrier type. However, for other systems, it is slightly lower for electrons compared to holes. This result was obtained experimentally for Ultrathin Black Phosphorus with Cu adatoms [70]. Furthermore, FIG. 4f illustrates that the lower the concentration of vacancies, the more pronounced the difference between the mobility of holes and that of electrons. The presence of a midgap state in the gap changes the value of $N_{e}$ (because defects here are neutral). It is therefore possible to compare mobility values, but it is more difficult to compare the values of $N_{e}$.
IV.2 Non-resonant scatterers
The Anderson disorder (non-resonant scatterers, see Sec. III) is randomly distributed on all layers of the different phosphorus systems studied. FIG. 5a,b and FIGs. S8 in SM [32] show the DOS for monolayer, 2, 3, 4, 5 layers, and bulk BP for different magnitudes $W$ of disorder. It is easy to see that the disorder decreases the gap for all systems (FIGs. 5c,d) unlike the case of resonant scatterers. For the bulk and high disorder ($W=4$), the gap becomes very small ($E_{g}\simeq 0.1$ eV). Most of the recent experiments show the effect of functionalization rather on the surface of BP multilayers to protect it from the deterioration coming from the ambient air. In the case when it is non-covalent, the band gap shrinks down monotonically by increasing the concentration of adsorbates in the topmost layer of few-layer BP.
@@ IV.2 Non-resonant scatterers
For large values of $W$, the gap gradually reduces and becomes zero, for bulk and multilayer systems, indicating a semi-conductor/semi-metal transition (see FIGs. S6 and S8 in SM [32]). This result is in good agreement with recent experiments for non-covalent functionalization  [71, 72, 25], which validates our hypothesis that the Anderson disorder can model non-covalent adsorbates.
@@
In SM [32], we present the effect of the Anderson disorder only in the top layer of few-layer BP on the band gap (FIG. S7). This shows that the effect of disorder on the gap is qualitatively similar to the case with Anderson disorder randomly distributed across all layers.
@@
FIGs. 5e,f present the armchair microscopic conductivity $\sigma_{xx}$ and zigzag microscopic conductivity $\sigma_{yy}$, versus energy for monolayer and for 3 layers. As expected due to anisotropy, $\sigma_{xx}$ is always larger than $\sigma_{yy}$ for all values of $E$. It is important to emphasize that the states filling the gap through disorder do not contribute significantly to the conduction. Thus, even though the gap depends quite strongly on the intensity of the disorder, the mobility gap is much less dependent on it.
@@
FIG. 6 presents the effect of Anderson disorder on the conductivity versus charge carriers $N_{e}$. The behavior of the conductivity is almost the same for the few-layer BP and the bulk with a slight difference for monolayer, for low values of $N_{e}$ (FIGs. 6a,b). We can also see that $\sigma_{xx}$ is higher for holes than for electrons. The comparison between $\sigma_{xx}$ and $\sigma_{yy}$ is showed FIGs. 6c,d. Like the case of resonant scatterers (vacancies), the armchair conductivity is larger than the zigzag conductivity and the ratio $\sigma_{xx}/\sigma_{yy}$ has the same order as the one obtained in the case of resonant scatterers.
[…]
V Discusions and conclusion
@@
Experimentally, phosphorene often contains a large number of local defects that are not always well characterized in the literature. These static defects may be due to the functionalization by adsorbed atoms or molecules, the effect of the substrate, or other structural effects. We aim to study the electronic structure and quantum transport in the presence of a high concentration of these defects. To do this, we considered two extreme types of generic local defects: resonant and non-resonant defects. We used the Rudenko tight binding (TB) model [27, 28] to analyze monolayer Black Phosphorene (BP), few-layer BP, and bulk BP. The TB model used is an effective four-band Hamiltonian, i.e. one effective orbital per P atom which is a $p_{z}$-like orbital. This makes it easy to simulate resonant adsorbates by vacancies (vacant P atoms) and non-resonant adsorbates by Anderson disorder (on-site energy disorder).
@@
Concerning the electronic structure, the two types of defect have very different effects on the gap. The resonant scatterers create midgap states that expand when the concentration $c$ of defects increases until the gap is filled. Whereas, non-resonant disorder reduces the gap value by filling the gap edges when its magnitude $W$ increases, As for resonant disorder, a high concentration of non-resonant disorder can fill the gap. This extreme case is easier to achieve for a multi-layer BP and the bulk BP which have a lower gap.
@@
Quantum diffusion calculations show that resonant and non-resonant defects act similarly on electrical conductivity. Indeed, even for the high defect concentrations we studied, the gap states of resonant defects and the gap edge states of non-resonant defects contribute very little to conductivity, and so the mobility gap remains close to the gap of the defect-free system. This suggests that electronic transport should be fairly well described by a semi-classical Bloch-Boltzamn-type approach, even when defects strongly modify the electronic structure. Particularly, the conductivity of states outside the gap is inversely proportional to the resonant defect concentration $c$. Note that for sufficiently small amounts of defect, when the gap is always present, it is possible to calculate electron mobility, and we obtain values that are entirely comparable to the experimental results.
@@
As expected, the strong anisotropy of BP’s atomic structure induces anisotropy in conductivity and mobility. This anisotropy is stronger in the case of the monolayer BP. And so from 2-layer BP, it is pretty close to that of the bulk BP. Our calculations show that this asymmetry persists in the presence of the two types of defect we have studied. And here again, it is difficult to find very different behavior between resonant defects and non-resonant defects. Comparing our calculations with the conductivity or mobility values found experimentally is difficult, as the defect concentration is not known. However, the anisotropy should be found experimentally. In particular, it would be possible to find experimentally that the ratio $\sigma_{xx}/\sigma_{yy}$ versus the charge carrier density $N_{e}$ exhibits a well-defined peak for small $N_{e}$.
@@ Acknowledgments
Numerical calculations have been performed at LPMC, Université de Tunis El Manar, and Centre de Calculs (CDC), CY Cergy Paris Université. We thank warmly Yann Costes and Baptiste Mary, CDC, for computing assistance. GTL and DM was supported by the ANR project J2D (ANR-15-CE24-0017). JJK thanks CY Advanced Studies for financial support.
Appendix A
@@
Electronic structure and conductivity in multilayer Black Phosphorene with resonant and non-resonant scatterers

### v2

@@ III Numerical method for transport
where $S$ is the surface and $E_{F}^{0}$ is the Fermi energy of the system without static defects, thus $E_{F}^{0}$ is at the middle of the gap.
Note that in the presence of defect band states in the gap that has a nonzero $\sigma^{\rm m}$, the DOS at $E_{F}$ can be non-zero, and therefore the mobility is not well defined numerically since the expression (8) may diverges for $N_{e}$ approaches zero. This is because the mobility is defined for a semiconductor but not for systems with states at $E_{F}$.
IV Results
In this section, we present the results on the effect of vacancies (resonant scatterers) and Anderson disorder (nonresonant scatterers) on the electronic properties of monolayer BP, few-layer BP, and bulk BP.
@@ IV.1 Resonant scatterers
Figure 3(a) illustrates the density of states (DOS) for a vacancy concentration of $c=1\%$ (resonant scatterers) relative to the total number of atoms in monolayer, few-layer BP, and bulk BP. Notably, there is a gap decrease as the number of layers increases. In line with defect-free behavior, the gap decreases as the number of layers increases.
@@
We can clearly distinguish the midgap state peak due to vacancies in the gap (Fig. 3(a)). It is not centered exactly at zero energy because the coupling terms beyond the first neighbors break the electron-hole symmetry of the bipartite Hamiltonian. Its width remains almost constant as the number of layers increases, showing that its energy dispersion is mainly due to intralayer couplings between midgap states. As the gap decreases with an increasing number of layers, the specific value of $c$ at which the midgap peaks fill the gap depends on the number of layers. For the monolayer, at the studied $c$ concentration, the gap is not filled, and we can distinguish secondary peaks (see Figs. S1 in the SM [36]). On the other hand, the gap is filled for $c>1-2\%$ in the case of 5 layers and the bulk, respectively (see Figs. S2 and S3 in the SM [36]). In Fig. 3(b), the points corresponding to these particular cases are not represented.
Figure 3(b) shows that overall, for few-layer BP (Fig. 3(d) for 3 layers) and bulk configurations, the gap increases as the vacancy concentration rises, except for the monolayer case where the gap remains nearly constant (Fig. 3(c)). That nonstandard behavior shows repulsion of continuum states by defect states (midgap states). From the point of view of a perturbative analysis, it is reasonable to think that this effect will be greater if the energies of the continuum states are closer to those of the midgap states, i.e. for a smaller gap corresponding to a larger number of layers. One way of understanding this phenomenon is perhaps to consider the fact that resonant defects distributed over a sub-lattice of a bipartite lattice can create a gap that increases with the concentration of defects [63, 74]. A detailed study would be needed to confirm this, but it is beyond the scope of this paper.
@@
Figures 3(e,f) show the microscopic conductivity multiplied by concentration, i.e. $c\sigma^{\rm m}$, versus energy for monolayer and trilayer according to the $x$-direction $\sigma^{\rm m}_{xx}$ (bold line) –armchair direction– and according to the $y$-direction $\sigma^{\rm m}_{yy}$ (dashed line) –zigzag direction–. See also Fig. S4 in SM [36] for bulk BP. The armchair conductivity $\sigma^{\rm m}_{xx}$ is greater than the zigzag conductivity $\sigma^{\rm m}_{yy}$ for all cases (monolayer, few-layer BP, and bulk). The proportionality between $\sigma^{\rm m}_{xx}$ and $\sigma^{\rm m}_{yy}$ remains the same for all systems. This ratio is a signature of the structural anisotropy of these materials, which results from an anisotropy of the effective masses of electrons and holes [10, 12, 7, 13].
Moreover, Figs. 3(e,f) show that the conductivity behaves according to the Bloch-Boltzmann approach, and then the conductivity at every $E$ is proportional to $1/c$. This result is consistent with experimental measurements of the monolayer BP grafted with PtCl2 groups [23]. It is also important to note that midgap states do not contribute much to conduction. Indeed, even if midgap states fill the gap, there is still a mobility gap. This may not be true for very high defect concentrations, which would allow strong conduction through defect states that would percolate into the structure to form a kind of midgap-state bands [74].
@@
We now study the effect of vacancies on the conductivity and mobility at room temperature ($T=300$ K) versus the charge carriers $N_{e}$ (Eqs. (6), (8) and (9)). The prediction of other works is not similar and depends roughly on the numerical study [35]. Figure 4(a) shows the armchair conductivity per layer versus $N_{e}$ for monolayer, few-layer BP, and bulk with a concentration of vacancies of $c=1\%$. The conductivity of holes is larger than that of electrons for high values of charge carriers. This difference is a direct consequence of the asymmetry of the DOS around the gap (see Sec. II comments on Fig. 2). In the cases of multilayer BP, the steeper DOS on the electron side indicates a lower velocity of charge carriers. It is therefore clear that the electron-hole asymmetry is much lower in the case of the monolayer for which there is almost no DOS asymmetry. The zigzag conductivity presents the same behavior (see Fig. S5 in SM [36]). These results are in agreement with the experimental results found by Xia et al.[7] for a BP thin film.
@@
To compare the effect of the concentration of vacancies, Fig. 4(b) presents the armchair conductivity versus $N_{e}$ for monolayer (bold line) and trilayer (dashed line) with different concentrations. As expected the conductivity decreases when $c$ increases, and the electrons/holes asymmetry is seen. In Fig. 4(c), the ratio $\sigma_{xx}/\sigma_{yy}$ is shown for monolayer, few-layer BP, and bulk versus $N_{e}$ for $c=1\%$. $\sigma_{xx}$ is more significant than $\sigma_{yy}$ regardless of the charge carrier nature. This anisotropy of BP is well known and has been highlighted theoretically [40, 75] and experimentally [41]. This is due to the anisotropy of the bands and therefore of the effective masses in the vicinity of the gap [10, 12, 7, 13]. It is easy to see that the anisotropy in the behavior of conductivity is more significant for electrons than for holes in all systems. However, the ratio $\sigma_{xx}/\sigma_{yy}$ is larger for the monolayer than for other systems in the case of holes, whereas in the case of electrons, this ratio becomes smaller compared to that of the multilayers. This ratio is also independent of the number of vacancies as found previously (Boltzmann transport). This is shown in Fig. 4(d) for monolayer and trilayer, and this is valid for all systems. Note that $\sigma_{xx}/\sigma_{yy}=\mu_{xx}/\mu_{yy}$.
@@
To enable simpler comparisons with experimental results, we plot the mobility at room temperature versus $N_{e}$ (Fig. 4(e)), according to formulas (8) and (9), for monolayer, few layers, and the bulk for $c=1\%$. In the case of the monolayer, the mobility is of the same order of magnitude regardless of the charge carrier type. However, for other systems, it is slightly lower for electrons compared to holes. This result was obtained experimentally for ultrathin black phosphorus with Cu adatoms [76]. Furthermore, Fig. 4(f) illustrates that the lower the concentration of vacancies, the more pronounced the difference between the mobility of holes and that of electrons. The presence of a midgap state in the gap changes the value of $N_{e}$ (because defects here are neutral). It is, therefore, possible to compare mobility values, but it is more difficult to compare the values of $N_{e}$.
IV.2 Nonresonant scatterers
@@
The Anderson disorder (nonresonant scatterers, see Sec. III) is randomly distributed across all layers of the different phosphorus systems studied. Figures 5(a,b) and Fig. S8 in the SM [36] show the DOS for monolayer, multilayers, and bulk BP for different magnitudes $W$ of disorder. It is easy to see that the disorder decreases the gap for all systems (Figs. 5(c,d)) unlike the case of resonant scatterers. This is expected for nonresonant defects, which generally create defect states at the band edges [73, 71]. For large values of $W$, the gap gradually reduces and becomes zero, for bulk ($E_{g}\simeq 0.1$ eV for $W=4$ eV) and multilayer systems, indicating a semiconductor/semimetal transition (see Figs. S6 and S8 in the SM [36]). This result is in good agreement with recent experiments for noncovalent functionalization [77, 78, 28], which validates our hypothesis that the Anderson disorder can model noncovalent adsorbates.
@@
In the SM [36], we show the effect of the Anderson disorder only in the top layer of a few-layer BP on the band gap (Fig. S7). This shows that the effect of disorder on the gap is qualitatively similar to the case with Anderson disorder randomly distributed across all layers.
@@
Figures 5(e,f) present the armchair microscopic conductivity $\sigma^{\rm m}_{xx}$ and zigzag microscopic conductivity $\sigma^{\rm m}_{yy}$, versus energy for monolayer and for 3 layers. As expected due to anisotropy, $\sigma^{\rm m}_{xx}$ is always larger than $\sigma^{\rm m}_{yy}$ for all values of $E$. It is important to emphasize that the states filling the gap through disorder do not contribute significantly to the conduction. Thus, even though the gap depends quite strongly on the intensity of the disorder, the mobility gap is much less dependent on it.
@@
Figure 6 presents the effect of Anderson disorder on the conductivity versus charge carriers $N_{e}$. The behavior of the conductivity is almost the same for the few-layer BP and the bulk with a slight difference for the monolayer, for low values of $N_{e}$ (Figs. 6(a,b)). We can also see that $\sigma_{xx}$ is higher for holes than for electrons. The comparison between $\sigma_{xx}$ and $\sigma_{yy}$ is shown in Figs. 6(c,d). Like the case of resonant scatterers (vacancies), the armchair conductivity is larger than the zigzag conductivity, and the ratio $\sigma_{xx}/\sigma_{yy}$ has the same order as the one obtained in the case of resonant scatterers.
In Figs. 6(e,f), the mobility at room temperature, calculated according to the formula (8), is shown. As for resonant scatterers, such a calculation can produce numerical artifacts for very small $N_{e}$ values. Indeed, when $N_{e}$ (Eq. (9)) tends to zero, the conductivity (Eq. (7)) must also tend to zero, which leads to a finite mobility $\mu$. But because of Gaussian Broadening in the calculation of $\sigma$ (Sec. III), $\sigma$ is overestimated and $\mu$ may be overestimated for $N_{e}\simeq 0$. However, this numerical artifact disappears, as soon as $N_{e}$ is not zero. Experimentally, it has been shown that for small solvent molecules, which are noncovalent scatterers, and at concentrations on the order of $10^{13}$ cm-2 and $10^{14}$ cm-2, the hole mobility is higher than that of electrons [79]. That is in good agreement with our results for low $N_{e}$ values.
V Discussions and conclusion
@@
Experimentally, phosphorene often contains a large number of local defects, which are not always well characterized in the literature. These static defects can be due to functionalization by adsorbed atoms or molecules, the effect of the substrate or other structural effects. We aim to study the electronic structure and quantum transport in the presence of a high concentration of these defects. To do this, we considered two extreme types of generic local defects: resonant and nonresonant defects. We have used the Rudenko TB model [30, 31] to analyze monolayer black phosphorene (BP), few-layer BP, and bulk BP. The TB model used is an effective four-band Hamiltonian, i.e. one effective orbital per P atom, which is a $p_{z}$-like orbital. This makes it easy to simulate resonant adsorbates by vacancies (vacant P atoms) and nonresonant adsorbates by Anderson disorder (on-site energy disorder).
@@
Concerning the electronic structure, the two types of defects have very different effects on the gap. The resonant scatterers create midgap states that expand when the concentration $c$ of defects increases until the gap is filled. In contrast, nonresonant disorder reduces the gap value by filling the gap edges when its magnitude $W$ increases. As in the case of resonant disorder, a high concentration of nonresonant disorder can fill the gap. This extreme case is easier to achieve for a multilayer BP and the bulk BP, which have a lower gap.
@@
Quantum diffusion calculations show that resonant and nonresonant defects act similarly on electrical conductivity. Indeed, even for the high defect concentrations we studied, the gap states of resonant defects and the gap edge states of nonresonant defects contribute very little to conductivity, and so the mobility gap remains close to the gap of the defect-free system. This suggests that electronic transport should be fairly well described by a semi-classical Bloch-Boltzmann-type approach, even when defects strongly modify the electronic structure. Particularly, the conductivity of states outside the gap is inversely proportional to the resonant defect concentration $c$. Note that for sufficiently small amounts of defects, when the gap is always present, it is possible to calculate the electron mobility, and we obtain values that are comparable to the experimental results.
@@
As expected, the strong anisotropy of BP’s atomic structure induces anisotropy in conductivity and mobility. This anisotropy is stronger in the case of monolayer BP. From a 2-layer BP, it is pretty close to that of the bulk BP. Our calculations show that this anisotropy persists in the presence of the two types of defects we have studied. And here again, it is difficult to find very different behavior between resonant defects and nonresonant defects. Comparing our calculations with the conductivity or mobility values found experimentally is difficult, as the defect concentration is not known. However, the anisotropy should be found experimentally. In particular, it would be possible to find experimentally that the ratio $\sigma_{xx}/\sigma_{yy}$ versus the charge carrier density $N_{e}$ exhibits a well-defined peak for small $N_{e}$. We also showed that the hole conductivity is greater than that of the electrons, particularly in multilayers and bulk, and this is valid for both types of defects. We can relate this effect to the asymmetry observed in their DOS without defects.
@@ Acknowledgments
Numerical calculations were conducted at LPMC, Université de Tunis El Manar, and Centre de Calculs (CDC), CY Cergy Paris Université. We warmly thank Yann Costes and Baptiste Mary from CDC for their computing assistance. GTL and DM were supported by the ANR projects J2D (ANR-15-CE24-0017) and FlatMoi (ANR-21-CE30-0029). JJK acknowledges financial support from CY Advanced Studies.
@@
Electronic structure and conductivity in functionalized multilayer black phosphorene